# The $k$-photon quantum Rabi model

### v1

@@ Abstract
A generalization of the quantum Rabi model is obtained by replacing the linear (dipole) coupling between the two-level system and the radiation mode by a non-linear expression in the creation and annihilation operators, corresponding to multi-photon excitations. If each spin flip involves $k$ photons, it is called the “$k$-photon” quantum Rabi model. While the formally symmetric Hamilton operator is self-adjoint in the case $k=2$, it is demonstrated here that the Hamiltonian is not self-adjoint for $k\geq 3$. Therefore it does not generate a unitary time evolution and is unphysical. This result cannot be obtained by numerical calculations in finite-dimensional spaces which attempt to approximate an unbounded operator by a finite-rank operator.
@@ 1 Introduction
Nevertheless, the model $H_{kp}$ has been studied from a theoretical and mathematical viewpoint for several years, especially the case $k=2$ duan , which features the “spectral collapse” phenomenon for sufficiently strong coupling, $g>g_{c}=\omega/2$. For values $g<g_{c}$ the model can be physically implemented via trapped ions felicetti . It was conjectured in ng-99 that $H_{2p}$ is no longer self-adjoint for $g>g_{c}$. The argument was based on the observation that for $g>\omega/2$ and $\Delta=0$, the Hamiltonian (2) can be mapped to the direct sum of an harmonic oscillator with inverted potential and its Fourier transform ng-99 . Therefore, it has no ground state and indeed no eigenvectors, the pure point spectrum is empty. However, the one-dimensional Schrödinger operator with potential $V=-|x|^{2}$ is known to be self-adjoint wienholtz ; hellwig ; kalf . The spectrum is continuous and spans the whole real axis, similar to the position operator $\hat{x}$ and the momentum operator $\hat{p}$. It follows then from the Kato-Rellich theorem reed-2 that $H_{2p}$ is self-adjoint for $\Delta\neq 0$, because $\Delta\sigma_{z}$ is bounded.
The case $k\geq 3$ has been studied in lo-98 and zhang2017 . Lo et al., as well as Zhang, come to the conclusion that the operator $H_{kp}$ is “ill-defined”. While Zhang only states that the pure point spectrum is empty, which does not exclude self-adjointness, it is correctly claimed in lo-98 that $H_{kp}$ is not self-adjoint for $k\geq 3$. To show this, Lo et al. use Nelson’s theorem in the reverse, i.e. it is assumed that an operator $\hat{A}$ is not self-adjoint if some dense set of vectors $\{\phi_{n}\}$ in the Hilbert space $\mathcal{H}$ yields no convergence of the expression
@@
In the present paper, it is proven with Bargmann space methods and asymptotic analysis that $H_{kp}$ has a continuous spectrum (in fact, all of $\mathbb{C}$), while the eigenvectors are normalizable  [^1] This seems to contradict Zhang’s statement in zhang2017 . But Zhang’s scalar product differs from (5), so it cannot be compared to the present calculation..
This result entails that $H_{kp}$ is not symmetric on its domain $\mathcal{D}(H_{kp})$ but only on a smaller set $\mathcal{D}^{\prime}$ (see section 3) and must act there trivially: either $\mathcal{D}^{\prime}\cap H_{kp}(\mathcal{D}^{\prime})=\emptyset$ or the single eigenvalue of $H_{kp}$ is zero.
At face value, $H_{kp}$ seems to be symmetric because it contains the sum of operators $\hat{a}^{k}$ and $\hat{a}^{\dagger k}$ which are mutually adjoint in $L^{2}(\mathbb{R})$. The total Hilbert space of $H_{kp}$ is $\mathbb{C}^{2}\otimes L^{2}(\mathbb{R})$. In the following, we shall study the problem in the Bargmann space of analytic functions $\phi(z)$, ${\cal B}$, which is isomorphic to $L^{2}(\mathbb{R})$ bargmann ; vourdas . In this space, $\hat{a}$ is given by the derivative ${\partial_{z}}$ and $\hat{a}^{\dagger}$ by multiplication with $z$. The operators $z$ and ${\partial_{z}}$ are mutually adjoint in ${\cal B}$ with the scalar product
@@ 2.1 $\bm{k=3}$
$\phi_{1}(t)\sim\exp\left(\frac{1}{2}r^{2}\pm\frac{\omega}{3}e^{-i3\theta/2}r\right)z(t)^{\rho}\big{(}c_{0}+\mathcal{O}(r^{-1})\big{)}.$
Because $\textrm{arg}(\gamma)\in\{\pi,\pm\pi/3\}$, the factor $\exp(\beta z)$ has unit modulus along the critical line and does not determine the finiteness of $|\phi_{1}|$. The convergence of the integral in $\langle\phi_{1}|\phi_{1}\rangle$ is then given by the value of $\rho(\gamma,\beta)$. The explicit calculation yields $\rho(\gamma,\beta)=-2$ for all possible $\gamma,\beta$. Because $\Re(\rho)<-1/2$, it follows that $|\phi_{1}|<\infty$. This is true whether a Stokes sector contains a critical line or not, thus knowledge of the Stokes multipliers is not needed, as proven in braak-23 . This means that the eigenvalue equation has normalizable solutions for all $E\in\mathbb{C}$. Therefore, the operator $H_{3p}$ is not self-adjoint nor has it self-adjoint extensions $H_{3p}^{ext}$ with $\mathcal{D}(H_{3p}^{ext})\subset{\cal B}$ because $\mathcal{D}(H_{3p})\subset{\cal B}$ already contains an uncountable set of (not mutually orthogonal) eigenfunctions with different eigenvalues.
@@ 2.2 $\bm{k=4}$
whose real part is always less than $-3/2$. We conclude again that all formal solutions of (13) are normalizable and $H_{4p}$ is therefore not self-adjoint.
@@ 3 The case $\bm{k\geq 5}$
Both exponents are less than $-1/2$ for $k>2$. Together with the results from sections 2.1 and 2.2, it follows that $H_{kp}$ is not self-adjoint in ${\cal B}$ for $k\geq 3$. Moreover, $H_{kp}$ is only “formally” symmetric on $\mathcal{D}(H_{kp})$: The expectation value of $z^{k}$,
@@
with certain positive constants $C_{1},C_{2}$ and $R$ large enough, is not finite for $\rho=\rho_{+}$, which means that the operator $H_{kp}$ is not symmetric on all of $\mathcal{D}(H_{kp})$ but only on a much smaller domain $\mathcal{D}^{\prime}$. If $H_{kp}$ is restricted to $\mathcal{D}^{\prime}$, it can have there no eigenvectors with non-zero eigenvalue.
@@ 4 Conclusions
We have shown that the $k$-photon quantum Rabi model is unphysical in the strict sense for $k\geq 3$, because its Hamilton operator $H_{kp}$ is, though apparently hermitian, not self-adjoint. Therefore, according to Stone’s theorem reed-1 , it cannot generate a unitary time development of quantum states. This feature of $H_{kp}$ is not detectable by any numerical evaluation which necessarily operates in a truncated, finite-dimensional Hilbert space. Any such calculation is misleading, even if a calculation of the spectrum seems to converge ng-99 . The employed method uses the analytically accessible asymptotics of formal solutions to the eigenvalue equation, although it contains arbitrarily high derivatives and is quite different from the commonly studied case where the ($d$-dimensional) Laplacian is augmented by some potential. Remarkably, the Stokes phenomenon, though certainly present, has no bearing on the conclusions. This is due to the vanishing of the exponent of second kind with order one for $k\geq 4$. If $\beta$ does not vanish, it may vary from Stokes sector to Stokes sector. Thus the usual lateral connection problem reappears which is hard to solve by analytic means. Finally, we note that the $k$-photon coupling in (2) may acquire physical meaning in the Dicke models with more than one qubit. Under certain conditions on the parameters of the Dicke models, there exist finite-dimensional invariant sub-spaces with bounded photon content, so-called “dark-like states”. The first such state was discovered numerically in chilingaryan2013 and mathematically identified as exact eigenstate in peng2014 . Later on, these states have been found in several generalizations of the Dicke model peng2015 and are highly relevant for applications in quantum information technology because they can be used for the fast generation of $W$-states peng2021 . For these states to exist, several qubits are required. If the qubit frequencies are different and fine-tuned to the mode frequency (but not to the coupling), a condition easily realizable in current exeprimental platforms, there are exact eigenstates of the Dicke Hamiltonian with finite photon number and energy which is independent from the coupling between qubits and radiation mode peng2015 . These states bear some resemblance to the quasi-exact states in the asymmetric quantum Rabi model reyes-bustos2022 and hint at another type of hidden symmetry. Very recently, there has been an attempt to identify this symmetry in the asymmetric generalization of the two-qubit Dicke model lei2023 . Regarding the $k$-photon model, there are also such exact eigenstates with finite photon number. If restricted to just these states, the Hamiltonian becomes clearly a finite-dimensional diagonal matrix with real entries and is trivially self-adjoint. But the analysis above has shown that no self-adjoint extension of the domain of $H_{kp}$ beyond these states exists.
@@ Acknowledgements.
I wish to thank Fumio Hiroshima for enlightening discussions. This work was funded by the Deutsche Forschungsgemeinschaft through grant no. 439943572.

### v2

@@ Abstract
A generalization of the quantum Rabi model is obtained by replacing the linear (dipole) coupling between the two-level system and the radiation mode by a non-linear expression in the creation and annihilation operators, corresponding to multi-photon excitations. If each spin flip involves $k$ photons, it is called the “$k$-photon” quantum Rabi model. While the formally symmetric Hamilton operator is self-adjoint in the case $k=2$, it is demonstrated here that the Hamiltonian is not self-adjoint for $k\geq 3$. Therefore it does not generate a unique unitary time evolution and is unphysical. This result cannot be obtained by numerical calculations in finite-dimensional spaces which attempt to approximate an unbounded operator by a finite-rank operator.
@@ 1 Introduction
Nevertheless, the model $H_{kp}$ has been studied from a theoretical and mathematical viewpoint for several years, especially the case $k=2$ duan , which features the “spectral collapse” phenomenon for sufficiently strong coupling, $g>g_{c}=\omega/2$. For values $g<g_{c}$ the model can be physically implemented via trapped ions felicetti . For the case $g>g_{c}$, it was conjectured in ng-99 that $H_{2p}$ is not self-adjoint: It is easily seen that for $g>\omega/2$ and $\Delta=0$, the Hamiltonian (2) can be mapped to a harmonic oscillator with inverted potential. Therefore, it has no ground state and indeed no eigenvectors, the pure point spectrum is empty. However, the harmonic oscillator with inverted potential is known to be self-adjoint wienholtz ; hellwig ; kalf and the continuous spectrum spans the whole real axis, similar to the position operator $\hat{x}$. The same applies to $H_{2p}$ braak-23 . The case $k\geq 3$ has been studied in lo-98 and zhang2017 . Lo et al., as well as Zhang, come to the conclusion that the operator $H_{kp}$ is “ill-defined”. While Zhang only states that the pure point spectrum is empty, which does not exclude self-adjointness, it is correctly claimed in lo-98 that $H_{kp}$ is not self-adjoint for $k\geq 3$. To show this, Lo et al. use Nelson’s theorem in the reverse, i.e. it is assumed that an operator $\hat{A}$ is not self-adjoint if some dense set of vectors $\{\phi_{n}\}$ in the Hilbert space $\mathcal{H}$ yields no convergence of the expression
@@
In the present paper, it is proven with Bargmann space methods and asymptotic analysis that $H_{kp}$ has a continuous spectrum (in fact, all of $\mathbb{C}$), while the eigenvectors are normalizable [^1] This seems to contradict Zhang’s statement in zhang2017 . But Zhang’s scalar product differs from (5), so it cannot be compared to the present calculation.. However, these elements of $L^{2}(\mathbb{R})$ are not located $\mathcal{D}(\hat{A})=\mathcal{D}(\hat{a}^{k})\cap\mathcal{D}(\hat{a}^{\dagger k})$.
$H_{kp}$ is formally symmetric because it contains the sum of operators $\hat{a}^{k}$ and $\hat{a}^{\dagger k}$ which are mutually adjoint in $L^{2}(\mathbb{R})$. The total Hilbert space of $H_{kp}$ is $\mathbb{C}^{2}\otimes L^{2}(\mathbb{R})$. In the following, we shall study the problem in the Bargmann space of analytic functions $\phi(z)$, ${\cal B}$, which is isomorphic to $L^{2}(\mathbb{R})$ bargmann ; vourdas . In this space, $\hat{a}$ is given by the derivative ${\partial_{z}}$ and $\hat{a}^{\dagger}$ by multiplication with $z$. The operators $z$ and ${\partial_{z}}$ are mutually adjoint in ${\cal B}$ with the scalar product
@@ 2.1 $\bm{k=3}$
$\phi_{1}(z)\sim\exp\left(\frac{1}{2}r^{2}\pm\frac{\omega}{3}e^{-i3\theta/2}r\right)z^{\rho}\big{(}c_{0}+\mathcal{O}(r^{-1})\big{)}.$
Because $\textrm{arg}(\gamma)\in\{\pi,\pm\pi/3\}$, the factor $\exp(\beta z)$ has unit modulus along the critical line and does not determine the finiteness of $|\phi_{1}|$. The convergence of the integral in $\langle\phi_{1}|\phi_{1}\rangle$ is then given by the value of $\rho(\gamma,\beta)$. The explicit calculation yields $\rho(\gamma,\beta)=-2$ for all possible $\gamma,\beta$. Because $\Re(\rho)<-1/2$, it follows that $|\phi_{1}|<\infty$. This is true whether a Stokes sector contains a critical line or not, thus knowledge of the Stokes multipliers is not needed, as proven in braak-23 . This means that the eigenvalue equation has normalizable solutions for all $E\in\mathbb{C}$ and all six combinations of $\gamma$ and $\beta$. This entails that $H_{3p}$ is not self-adjoint, although it may have self-adjoint extensions because the defect indices are $\{6,6\}$ reed-2 . However, these extensions are not unique and it is not possible to define the extended domain of $H_{3p}$ by restricting the behavior of $\phi_{1}(z)$ for $z\rightarrow\infty$ in a physically meaningful way. For example, just demanding a finite photon content, $\langle\phi_{1}|\hat{a}^{\dagger}\hat{a}|\phi_{1}\rangle\stackrel{{\scriptstyle!}}{{<}}\infty$, does not remove enough of the uncountably many, not mutually orthogonal eigenstates.
@@ 2.2 $\bm{k=4}$
whose real part is always less than $-3/2$. We conclude again that all formal solutions of (13) are normalizable and $H_{4p}$ is therefore not self-adjoint. Moreover, the argument of section 2.1 about the feasability of possible self-adjoint extensions applies here as well: The Hamiltonian $H_{4p}$ does not describe a physically realizable system.
@@ 3 The case $\bm{k\geq 5}$
Both exponents are less than $-1/2$ for $k>2$. Together with the results from sections 2.1 and 2.2, it follows that $H_{kp}$ is not self-adjoint in ${\cal B}$ for $k\geq 3$. The constructed solutions of the eigenvalue equation are elements of $\mathcal{H}=\mathcal{B}\otimes\mathbb{C}^{2}$, but not of $\mathcal{D}(H_{kp})$, as seen as follows. The expectation value of $z^{k}=\hat{a}^{\dagger k}$,
@@
with certain positive constants $C_{1},C_{2}$ and $R$ large enough, is not finite for $\rho=\rho_{+}$ and as $\mathcal{D}(H_{kp})\subset\mathcal{D}(\hat{a}^{\dagger k})$, the eigenfunctions must be located in some (not unique) self-adjoint extension of $H_{kp}$.
@@ 4 Conclusions
We have shown that the $k$-photon quantum Rabi model is unphysical in the strict sense for $k\geq 3$, because its Hamilton operator $H_{kp}$ is, though apparently hermitian, not self-adjoint. Therefore, according to Stone’s theorem reed-1 , it cannot generate a unitary time development of quantum states. This feature of $H_{kp}$ is not detectable by any numerical evaluation which necessarily operates in a truncated, finite-dimensional Hilbert space. Any such calculation is misleading, even if a calculation of the spectrum seems to converge ng-99 . The employed method makes use the analytically accessible asymptotics of formal solutions to the eigenvalue equation, although it contains arbitrarily high derivatives and is quite different from the commonly studied case where the ($d$-dimensional) Laplacian is augmented by some potential. Remarkably, the Stokes phenomenon, though certainly present, has no bearing on the conclusions. This is due to the vanishing of the exponent of second kind with order one for $k\geq 4$. If $\beta$ does not vanish, it may vary from Stokes sector to Stokes sector. Thus the usual lateral connection problem reappears which is hard to solve by analytic means. Finally, we note that the $k$-photon coupling in (2) may acquire physical meaning in the Dicke models with more than one qubit. Under certain conditions on the parameters of the Dicke models, there exist finite-dimensional invariant sub-spaces with bounded photon content, so-called “dark-like states”. The first such state was discovered numerically in chilingaryan2013 and mathematically identified as exact eigenstate in peng2014 . Later on, these states have been found in several generalizations of the Dicke model peng2015 and are highly relevant for applications in quantum information technology because they can be used for the fast generation of $W$-states peng2021 . For these states to exist, several qubits are required. If the qubit frequencies are different and fine-tuned to the mode frequency (but not to the coupling), a condition easily realizable in current exeprimental platforms, there are exact eigenstates of the Dicke Hamiltonian with finite photon number and energy which is independent from the coupling between qubits and the radiation mode peng2015 . These states bear some resemblance to the quasi-exact states in the asymmetric quantum Rabi model reyes-bustos2022 and hint at another type of hidden symmetry. Very recently, there has been an attempt to identify this symmetry in the asymmetric generalization of the two-qubit Dicke model lei2023 . The $k$-photon Dicke model possesses also such exact eigenstates with finite photon number. If restricted to just these states, the Hamiltonian becomes clearly a finite-dimensional diagonal matrix with real entries and is trivially self-adjoint. But the analysis above demonstrates that the generic eigenstates of a system with $k$-photon coupling cannot be meaningfully associated with any self-adjoint extension of its Hamilton operator.
@@ Acknowledgements.
The author wishes to thank for enlightening discussions with Fumio Hiroshima, Daniel Burgarth and Davide Lonigro. This work was funded by the Deutsche Forschungsgemeinschaft through grant no. 439943572.